OpenStudy (anonymous):

Solve equation for y. then find the value of y for each value of x 5x = -4y + 4; x = 1, 2, 3

OpenStudy (anonymous):

5x = -4y + 4 4y=-5x+4 y=(-5/4)x +1 when x = 1, 2, 3 y=-1/4, -3/2,-11/4